# Uniform irreducibility of Galois action on the $\ell$-primary part of Abelian $3$-folds of Picard type

### v1

@@ 3.4. The Bombieri–Lang Conjecture for Picard modular surfaces
In this part we follow the general strategy of [10] by adapting it to the case where the level is not neat, the main point being to show that the irregularity of the Picard modular surfaces under consideration is at least $3$. This requires different techniques also providing a new proof to the cases treated in loc. cit..
@@
Consider an automorphic representation $\pi\in\Pi(\lambda_{c}\chi_{M})$ having non-zero $K^{\prime\prime}_{0}(\ell^{r})$-invariants. By Propositions 1.4 and 1.6, for all finite place $v\neq\ell$, we have $\pi_{v}=\pi_{n,v}$ with $\chi_{v}$ unramified, and conversely it follows from (7) that the line $\bigotimes_{p\mid D}\pi_{n,p}^{K^{\prime\prime}_{p}}$ is fixed by $K^{\prime\prime}_{D}$. By (14) that $\pi$ contributes to $q(Y_{\Gamma^{\prime\prime}_{0}(\ell^{r})}^{\ast})$ only when $\pi_{\ell}=\pi_{n,\ell}$ and $W(\lambda_{c}^{3}\chi_{M}^{3})=1$, or $\pi_{\ell}=\pi_{c,\ell}$ and $W(\lambda_{c}^{3}\chi_{M}^{3})=-1$. As the choice of a character $\chi_{\ell}:\mathbf{Z}_{\ell^{2}}^{1}\to\mathbf{C}^{\times}$ uniquely determines, up to a class character, a global character $\chi_{M}$ unramified outside $\ell$, the irregularity formula becomes:
(17) $\tfrac{1}{h}\cdot q\left(Y_{\Gamma^{\prime\prime}_{0}(\ell^{r})}^{\ast}\right)=\sum_{W(\lambda_{c}^{3}\chi_{M}^{3})=-1}\dim(\pi_{c,\ell}^{I_{r}})+\sum_{W(\lambda_{c}^{3}\chi_{M}^{3})=1}\dim(\pi_{n,\ell}^{I_{r}}),$
and it can be rendered even more explicit using Lemma 3.3.
The study of the conductor of the super-cuspidal non-generic representation $\pi_{c,\ell}$ appears to be very delicate even in the depth 0 case (i.e. trivial $\chi$), where some preliminary computations suggest that it does not have $K_{T}$-invariants. Consequently we will use the lower bound on the irregularity corresponding to the contribution of $\pi$ which are the everywhere non-tempered.
@@
We recall that $D>3$ is odd and let $h$ be the class number of $M=\mathbf{Q}(\sqrt{-D})$.
@@
    If $D\equiv 7\pmod{8}$ then $q\left(Y^{\ast}_{\Gamma_{0}^{\prime\prime}(27)}\right)\geqslant h$ and $q\left(Y^{\ast}_{\Gamma_{0}^{\prime\prime}(\ell^{3})}\right)\geqslant 3h$, for $\ell\geqslant 7$.
@@
The proof is based on the inequality $\left(Y_{\Gamma^{\prime\prime}_{0}(\ell^{r})}^{\ast}\right)=h\cdot\sum_{W(\lambda_{c}^{3}\chi_{M}^{3})=1}\dim(\pi_{n,\ell}^{I_{r}})$ resulting from (17) and the results on root numbers in §$3.$
(i) As $W(\lambda_{c}^{3})=1$ we can take $\chi=\mathbf{1}$. The claim follows from Proposition 1.6 and Corollary 1.9.
(ii) As $W(\lambda_{c}^{3})=-1$ we use here a tamely ramified $\chi_{\ell}$ to switch the sign. By Lemma 3.3, in order to have $W(\lambda_{c}^{3}\chi_{M}^{3})=-W(\lambda_{c}^{3})=1$, we need $\chi_{\ell}^{3}$ to be non-trivial and $\chi_{\ell}(-1)=1$ For $\ell\geqslant 3$, there are precisely $\frac{\ell-1}{2}$ choices for $\chi_{\ell}$, if $3\nmid(\ell+1)$, and $\frac{\ell-5}{2}$ choices, if $3\mid(\ell+1)$. In particular, there are at least $3$ choices for all $\ell\geqslant 7$. ∎
The following results from Faltings [12] (see [10, §3.2] for detail).
@@
Let $\widetilde{K}$ an open compact subgroup of $\widetilde{G}(\mathbf{A}_{f})$ and let $\Gamma=\widetilde{G}(\mathbf{Q})\cap\widetilde{K}\widetilde{G}(\mathbf{R})$. If $q\left(Y^{\ast}_{\Gamma}\right)\geqslant 3$, then $Y^{\ast}_{\widetilde{K}}$ satisfies the Bombieri–Lang Conjecture, i.e., any number field $k$ the $k$-rational points of $Y^{\ast}_{\widetilde{K}}$ are not Zariski dense.
@@
If $D\equiv 3\pmod{8}$, then the Bombieri–Lang Conjecture holds for $Y_{\widetilde{K}^{\prime\prime}(\ell^{4})}$, and even for $Y^{\ast}_{\widetilde{K}^{\prime\prime}}$, when $h\geqslant 3$. If $D\equiv 7\pmod{8}$, then the Bombieri–Lang Conjecture holds for $Y^{\ast}_{\widetilde{K}^{\prime\prime}(\ell^{3})}$ for $\ell\geqslant 7$, and for $Y^{\ast}_{\widetilde{K}^{\prime\prime}(3^{7})}$.
As $2$ splits in $M$ for $D\equiv 7\pmod{8}$, we only exclude $\ell=5$ when $D\equiv 7\text{ or }23\pmod{40}$.
4. Uniform irreducibility of Galois images
4.1. Complex reflexions, elliptic elements and abelian families
[…]
Recall that non-scalar element $\gamma$ of the discrete subgroup $\Gamma\subset\widetilde{G}(\mathbf{R})$ (i.e. a non-trivial element in $\overline{\Gamma}=\Gamma/(\Gamma\cap M^{\times})$) has a fixed point in $\mathcal{H}$ if, and only if, $\gamma$ has finite order (this is because the stabilizers in $\widetilde{G}(\mathbf{R})$ of points in $\mathcal{H}$ are maximal compact subgroups). Such a $\gamma$ is called elliptic if it only fixes an isolated point in $\mathcal{H}$, otherwise it is called a complex reflexions.
Lemma 4.1.
Proof.
[…]
Lemma 4.2.
The set of fixed points in $Y^{\prime\prime}(\mathbf{C})$ is finite and defined over a finite extension of $M$.
Proof.
One proceeds by explicitly determining the elliptic elements of $\Gamma^{\prime\prime}$. ∎
After possibly enlarging $k$ we may assume that the set $Z$ from Lemma 4.2 is defined over $k$ and we denote by $U$ the complementary open in $Y^{\prime\prime}$. Following the discussion in §2.2 the open $U$ lifts to an open in the corresponding algebraic moduli stack, as such it is naturally endowed with a family $f:A\to U$ of abelian $3$-folds of Picard type.
@@ 4.2. Proof of Main Theorem
At different stages of the proof we will remove finite sent and deal with them in the last step. By §2.1 and §2.4 an abelian variety $A$ as in the Theorem B defines a $k$-rational point on $Y_{\widetilde{K}^{\prime\prime}}$, where $\widetilde{K}^{\prime\prime}$ is defined in (6). If $A[\ell^{r}]$ admits a full $k$-rational flag (or equivalently, a $k$-rational isotropic line), we claim that $A$ defines a $k$-rational point on $Y_{\widetilde{K}^{\prime\prime}_{0}(\ell^{r})}$, where $\widetilde{K}^{\prime\prime}_{0}(\ell^{r})\subset\widetilde{K}^{\prime\prime}$ is the subgroup whose component at $\ell$ consists of elements whose reduction modulo $\ell^{r}$ belong to the standard Borel subgroup $\widetilde{B}(\mathbf{Z}/\ell^{r}\mathbf{Z})$ of $\widetilde{G}(\mathbf{Z}/\ell^{r}\mathbf{Z})$. Indeed, by assumption one knows that there is some Borel subgroup containing the image of $\operatorname{Gal}_{k}$ acting on $A[\ell^{r}]$. However $\widetilde{G}(\mathbf{Q}_{\ell})$ acts transitively on isotropic lines (because isometries between hermitian subspaces always extend), hence all Borel subgroups are conjugated by $\widetilde{G}(\mathbf{Q}_{\ell})$, and in fact by $\widetilde{K}_{\ell}^{\circ}=\widetilde{G}(\mathbf{Z}_{\ell})$ (using Iwasawa decomposition). As $\widetilde{K}^{\prime\prime}$ is a normal subgroup of $\widetilde{K}^{\circ}$ we deduce that the Galois image is contained in $\widetilde{K}^{\prime\prime}\cap\widetilde{K}_{0}(\ell^{r})=\widetilde{K}^{\prime\prime}_{0}(\ell^{r})$, proving the claim.
By Corollary 3.6, $Y^{\ast}_{\widetilde{K}^{\prime\prime}_{0}(\ell^{4})}$ satisfies the Strong Bombieri–Lang Conjecture. In particular all its $k$-rational points lie in a subvariety $Z$ defined over $k$ which is a finite union of points and curves.
@@
Consider image $\Gamma_{C}\subset\widetilde{K}_{\ell}^{\circ}=\widetilde{G}(\mathbf{Z}_{\ell})$ of the étale fundamental group acting on the $\ell$-adic Tate module of the generic fiber of the family. By Cartan’s theorem (see [30, LG5.42]), $\Gamma_{C}$ is an $\ell$-adic Lie group hence admits a Lie algebra $\mathfrak{h}$. By Bogomolov [2] the Lie algebra $\mathfrak{h}_{\mathbf{Q}_{\ell}}$ is the Lie algebra of the Zariski closure of $\Gamma_{C}$ in $\widetilde{G}(\mathbf{Q}_{\ell})$, the latter being furthermore reductive over $\mathbf{Q}_{\ell}$ by Faltings [11, Thm. 3]. By the Mumford–Tate Conjecture, which is known for is known for abelian $3$-folds (see e.g. [5]), we know that $\mathfrak{h}_{\mathbf{Q}_{\ell}}$ is the Lie algebra of the Mumford–Tate group $\operatorname{MT}(A)\otimes\mathbf{Q}_{\ell}$. As $C$ has positive dimension, it has to contain non-CM points, whose Mumford–Tate group is not abelian. By Lemma 2.1, the Lie subalgebra $\mathfrak{h}_{\mathbf{Q}_{\ell}}\subset\mathfrak{su}(3,\mathbf{Q}_{\ell})$ contains a form of $\mathfrak{sl}(2,\mathbf{Q}_{\ell})$. By [4, Thm. 1.1] applied to the abelian family $f:A\to C$ there exist $B>0$ such that for all $x\in C(k)$ outside a finite set $C_{\rho}$ we have
@@
There exists $r=r(C)\in\mathbf{Z}$ such that $\Gamma_{x}\supset\exp\left(\mathfrak{su}(2,\ell^{r}\mathbf{Z}_{\ell})\right)$ for all $x\in C(k)\setminus C_{\rho}$.
@@
It follows that $r\geqslant r(C)$ and for all $A$ as above corresponding to a $k$-rational point on $C\setminus C_{\rho}$, $A[\ell^{r}]$ does not admit a full $k$-rational flag. Finally, a direct application of Theorem 2.3 to the finitely many remaining $k$-rational points yields an integer $r$ such that all $k$-rational points in $Y_{\widetilde{K}^{\prime\prime}_{0}(\ell^{r})}$ are of CM type, proving the Theorem.
4.3. Odd and ends
In a previous paper, we mistakenly switched $\pi_{c}$ and $\pi_{2}$ (which are in the same $L$-packet) in the proof of Proposition 3.8 in [10] which also had consequences for the assertions of Theorem 0.3 in loc. cit.. Providing $I_{r}$-invariants for its supercuspidal member $\pi_{c}$ seem to be a thorny issue. Even in the simplest case of $\pi_{c}\in\Pi(\lambda_{0})$ which can be shown to be of depth 0, i.e. coming from representation theory of finite groups, preliminary computation via Deligne–Lusztig theory suggest that it does not have $K_{T}$-invariants. This leaves us with no choice but to use $\pi_{n}$.
The methods of this paper based on ramified $\pi_{n}$ allow us to find a new proof of part (ii) of Theorem 0.3 in loc. cit., in fact yielding an even stronger result. The case (i) will be discussed elsewhere.

### v2

@@ 3.4. The Bombieri–Lang Conjecture for Picard modular surfaces
In this part we follow the general strategy of [11] by adapting it to the case where the level is not neat, the main point being to show that the irregularity of the Picard modular surfaces under consideration is at least $3$. This requires some new techniques, also providing new proofs to the cases treated in loc. cit.
@@
Consider an automorphic representation $\pi\in\Pi\left(\lambda_{c}\chi_{M},{\lambda^{-1}_{c}}_{\!\!\!\!\!|M^{1}}\chi^{-2}\omega\right)$ having non-zero $K^{\prime\prime}_{0}(\ell^{r})$-invariants. By Propositions 1.3 and 1.5, for all finites place $v\neq\ell$ we must have $\pi_{v}=\pi_{n,v}$ with $\chi_{v}$ unramified, in which case it follows from (14) that the line $\bigotimes\limits_{p\mid D}\pi_{n,p}^{K^{\prime\prime}_{p}}$ is fixed by $K^{\prime\prime}_{D}$. By (18), $\pi$ contributes to $q\left(Y_{\Gamma^{\prime\prime}_{0}(\ell^{r})}^{\ast}\right)$ if either $\pi_{\ell}=\pi_{n,\ell}$ and $W(\lambda_{c}^{3}\chi^{3}_{M})=1$, or $\pi_{\ell}=\pi_{c,\ell}$ and $W(\lambda_{c}^{3}\chi^{3}_{M})=-1$. As the choice of a character $\chi_{\ell}:\mathbf{Q}_{\ell^{2}}^{1}\to\mathbf{C}^{\times}$ determines uniquely, up to an ideal class character, a global character $\chi_{M}$ unramified outside $\ell$, the irregularity formula (18) becomes (we have omitted $\omega\in\widehat{\mathscr{C}\!\ell_{M}^{1}}$ as it is trivial on $\mathbf{Q}_{\ell^{2}}^{1}$)
(22) $\tfrac{1}{h}\cdot q\left(Y_{\Gamma^{\prime\prime}_{0}(\ell^{r})}^{\ast}\right)=\sum_{\chi_{\ell},\text{ s.t.}W(\lambda_{c}^{3}\chi_{M}^{3})=1}\dim(\pi_{n,\chi_{\ell}}^{I_{r}})+\sum_{\chi_{\ell},\text{ s.t.}W(\lambda_{c}^{3}\chi_{M}^{3})=-1}\dim(\pi_{c,\chi_{\ell}}^{I_{r}}),$
where $\Pi(\lambda_{0}\chi_{M,\ell},\chi^{-2}_{\ell})=\{\pi_{n,\chi_{\ell}},\pi_{c,\chi_{\ell}}\}$ denotes the local Arthur packet attached to the quadratic unramified character $\lambda_{0}:\mathbf{Q}_{\ell^{2}}^{\times}\to\mathbf{C}^{\times}$ and to $\chi_{\ell}:\mathbf{Q}_{\ell^{2}}^{1}\to\mathbf{C}^{\times}$.
Remark 3.4.
The study of the conductor of the supercuspidal non-generic representation $\pi_{c,\chi_{\ell}}$ appears to be very delicate. Indeed, even in the depth 0 case (i.e., trivial $\chi$), some preliminary computations suggest that it does not contain non-zero $K_{T}$-invariant vectors. In a previous paper ([11]), we mistakenly switched $\pi_{c}$ and $\pi_{2}$ (which are in the same $L$-packet) in the proof of Propositions 3.6 and 3.8 in loc. cit.. It did not affect Theorem 0.2 there at all, but in Theorem 0.1 when $W(\lambda^{3})=(-1)^{d}$, we need to replace $\Gamma_{1}(\mathfrak{C})$ by $\Gamma_{1}(\mathfrak{Cq})$ for any $\mathfrak{q}$ prime of $F$ relatively prime to $\mathfrak{C}$ and non-split in $M$. We overcome the difficulty caused by $\pi_{\mathfrak{q},c}$ not having $\Gamma_{1}(\mathfrak{q})$-invariants by using an appropriate twist $\lambda_{c}\chi_{M}$ of $\lambda_{c}$ for which $W(\lambda_{c}^{3}\chi^{3}_{M})=(-1)^{d-1}$ and the corresponding $\pi_{\mathfrak{q},n}$ has the requisite invariants. Adapting this twisting method at ramified places, allows us to establish a variant of Theorem 0.3 from loc. cit.. This will be taken up elsewhere.
Henceforth, we will only use a lower bound for the irregularity corresponding to the contribution of everywhere non-tempered automorphic representations. Namely, combining (22) with Lemma 3.3, we have
(23) $q\left(Y_{\Gamma^{\prime\prime}_{0}(\ell^{r})}^{\ast}\right)\geqslant h\cdot\sum_{\chi_{\ell}}\dim(\pi_{n,\chi_{\ell}}^{I_{r}}),$
where the sums runs over characters $\chi_{\ell}:\mathbf{Q}_{\ell^{2}}^{1}\to\mathbf{C}^{\times}$ such that $\chi_{\ell}(-1)=(-1)^{a(\chi_{\ell})+(D-3)/4}$.
@@
We recall that $D>3$ is odd and $h$ denotes the class number of $M=\mathbf{Q}(\sqrt{-D})$.
@@
    If $D\equiv 7\pmod{8}$ then $q\left(Y^{\ast}_{\Gamma_{0}^{\prime\prime}(27)}\right)\geqslant h$, $q\left(Y^{\ast}_{\Gamma_{0}^{\prime\prime}(5^{6})}\right)\geqslant 3h$ and $q\left(Y^{\ast}_{\Gamma_{0}^{\prime\prime}(\ell^{3})}\right)\geqslant 3h$, for $\ell\geqslant 7$.
@@
(i) We can take $\chi=\mathbf{1}$ in (23). The claim follows from Proposition 1.5 and Corollary 1.8.
(ii) In this case, we use tamely ramified $\chi_{\ell}$ in (23), i.e. $a(\chi_{\ell})=1$, such that $\chi_{\ell}(-1)=1$. For $\ell\geqslant 3$, there are precisely $\frac{\ell-1}{2}$ such characters, if $3\nmid(\ell+1)$, and $\frac{\ell-5}{2}$ choices, if $3\mid(\ell+1)$. In particular, there are at least $3$ choices for all $\ell\geqslant 7$. As $K_{T}$ contains a conjugate $I_{3}$, the claim then follows from Theorem 1.9 supplemented, when $\ell=5$, by a numerical computation showing that $\pi_{n,\chi_{5}}$ has non-zero $I_{6}$-invariants for $\chi_{5}$ such that $\chi_{5}(-1)=-1$ and $a(\chi_{5})=2$. ∎
The following results from Faltings [13] (see [11, §3.2] for details).
@@
Let $\widetilde{K}$ an open compact subgroup of $\widetilde{G}(\mathbf{A}_{f})$ and let $\Gamma=\widetilde{G}(\mathbf{Q})\cap\widetilde{K}\widetilde{G}(\mathbf{R})$. If $q\left(Y^{\ast}_{\Gamma}\right)\geqslant 3$, then $Y^{\ast}_{\widetilde{K}}$ satisfies the Bombieri–Lang Conjecture.
As $2$ splits in $M$ for $D\equiv 7\pmod{8}$, Proposition 3.5 has the following consequence.
@@
If $D\equiv 3\pmod{8}$, then the Bombieri–Lang Conjecture holds for $Y^{\ast}_{\widetilde{K}^{\prime\prime}(\ell^{4})}$, and even for $Y^{\ast}_{\widetilde{K}^{\prime\prime}}$, when $h\geqslant 3$. If $D\equiv 7\pmod{8}$, then the Bombieri–Lang Conjecture holds for $Y^{\ast}_{\widetilde{K}^{\prime\prime}(\ell^{3})}$ for $\ell\geqslant 7$, and also for $Y^{\ast}_{\widetilde{K}^{\prime\prime}(3^{7})}$ and $Y^{\ast}_{\widetilde{K}^{\prime\prime}(5^{6})}$.
@@ 3.5. Proof of Main Theorem
At different stages of the proof we will remove finite sets and deal with them in the last step. By §2.1 an abelian variety $A$ as in the Theorem B defines a $k$-rational point on $Y_{\widetilde{K}^{\prime\prime}}$, where $\widetilde{K}^{\prime\prime}$ is defined in (13). If $A[\ell^{r}]$ admits a full $k$-rational flag (or equivalently, a $k$-rational isotropic line), we claim that $A$ defines a $k$-rational point on $Y_{\widetilde{K}^{\prime\prime}_{0}(\ell^{r})}$, where $\widetilde{K}^{\prime\prime}_{0}(\ell^{r})\subset\widetilde{K}^{\prime\prime}$ is the subgroup whose component at $\ell$ consists of elements whose reduction modulo $\ell^{r}$ belong to the standard Borel subgroup $\widetilde{B}(\mathbf{Z}/\ell^{r}\mathbf{Z})$ of $\widetilde{G}(\mathbf{Z}/\ell^{r}\mathbf{Z})$. Indeed, by assumption one knows that there is some Borel subgroup containing the image of $\operatorname{Gal}_{k}$ acting on $A[\ell^{r}]$. However $\widetilde{G}(\mathbf{Q}_{\ell})$ acts transitively on isotropic lines (because isometries between hermitian subspaces always extend), hence all Borel subgroups are conjugated by $\widetilde{G}(\mathbf{Q}_{\ell})$, and in fact by $\widetilde{K}_{\ell}^{\circ}=\widetilde{G}(\mathbf{Z}_{\ell})$ (using Iwasawa decomposition). As $\widetilde{K}^{\prime\prime}$ is a normal subgroup of $\widetilde{K}^{\circ}$ we deduce that the Galois image is contained in $\widetilde{K}^{\prime\prime}\cap\widetilde{K}_{0}(\ell^{r})=\widetilde{K}^{\prime\prime}_{0}(\ell^{r})$, proving the claim.
By Corollary 3.7, $Y^{\ast}_{\widetilde{K}^{\prime\prime}_{0}(\ell^{7})}$ satisfies the Strong Bombieri–Lang Conjecture. In particular all its $k$-rational points lie in a subvariety $Z$ defined over $k$ which is a finite union of points and curves.
@@
As in §2.3 let $\Gamma_{C}\subset\widetilde{K}_{\ell}^{\circ}=\widetilde{G}(\mathbf{Z}_{\ell})$ be the image of the étale fundamental group acting on the $\ell$-adic Tate module of the generic fiber of the family. By Cartan’s theorem (see [30, LG5.42]), $\Gamma_{C}$ is an $\ell$-adic Lie group hence admits a Lie algebra $\mathfrak{g}_{\mathbf{Z}_{\ell}}$. By Bogomolov [2] the Lie algebra $\mathfrak{g}_{\mathbf{Q}_{\ell}}=\mathfrak{g}_{\mathbf{Z}_{\ell}}\otimes_{\mathbf{Z}_{\ell}}\mathbf{Q}_{\ell}$ is algebraic, namely it is the Lie algebra of the Zariski closure of $\Gamma_{C}$ in $\widetilde{G}(\mathbf{Q}_{\ell})$, the latter being furthermore reductive over $\mathbf{Q}_{\ell}$ by Faltings [12, Thm. 3]. By the Mumford–Tate Conjecture, which is known for is known for abelian $3$-folds (see e.g. [5]), we know that $\mathfrak{g}_{\mathbf{Q}_{\ell}}$ is the Lie algebra of the Mumford–Tate group $\operatorname{MT}(A)\otimes_{\mathbf{Q}}\mathbf{Q}_{\ell}$. As $C$ has positive dimension, it has to contain non-CM points, whose Mumford–Tate group is not abelian. By Lemma 2.3, the Lie subalgebra $\mathfrak{g}_{\mathbf{Q}_{\ell}}\cap\mathfrak{su}(3,\mathbf{Q}_{\ell})$ contains a form of $\mathfrak{sl}(2,\mathbf{Q}_{\ell})$. By [4, Thm. 1.1] applied to the abelian family $f:A\to C$ there exist $B>0$ such that for all $x\in C(k)$ outside a finite set $C_{\rho}$ we have
@@
There exists $r=r(C)\in\mathbf{Z}$ such that $\Gamma_{x}\supset\exp\left(\mathfrak{su}(2,\ell^{r}\mathbf{Z}_{\ell})\right)$ for all $x\in C(k){\smallsetminus}C_{\rho}$.
@@
It follows that for all $r\geqslant r(C)$ and for all $A$ as above corresponding to a $k$-rational point on $C{\smallsetminus}C_{\rho}$, $A[\ell^{r}]$ does not admit a full $k$-rational flag. Finally, applying Proposition 2.4 to the finitely many remaining $k$-rational points, yields an integer $r$ such that all $k$-rational points in $Y_{\widetilde{K}^{\prime\prime}_{0}(\ell^{r})}$ are of CM type, completing the proof of the Theorem.